# The Fundamental Parameters of Astrophysical Plasma Turbulence and its Dissipation: Nonrelativistic Limit

### v1

@@ 4 Review of Previous Turbulent Heating Models
The Howes 2010 (hereafter H10) turbulent heating model (Howes, 2010) was based on a turbulent cascade model (Howes et al., 2008a) that extended the GS95 turbulence scaling from MHD scales at $k_{\perp}\rho_{i}\lesssim 1$ into the kinetic regime at $k_{\perp}\rho_{i}\gtrsim 1$, generalizing the conjecture of critical balance to account for the dispersive nature of kinetic Alfvén waves at these sub-ion length scales (Biskmap et al., 1999; Cho & Lazarian, 2004; Krishan & Mahajan, 2004; Shaikh & Zank, 2005; Schekochihin et al., 2009). The H10 model was based on three primary assumptions: (i) the Kolmogorov hypothesis that the energy cascade is determined by local interactions (Kolmogorov, 1941); (ii) the conjecture that turbulence maintains a state of critical balance at all scales (Goldreich & Sridhar, 1995); and (iii) the speculation that linear collisionless damping rates are applicable even in the presence of the strong nonlinear interactions that mediate the turbulent cascade to small scales. The casade model used by H10 balanced at each perpendicular wavenumber $k_{\perp}\rho_{i}$ the nonlinear energy cascade rate given by the extended turbulent scaling theories with the linear collisionless damping rates at that value of $k_{\perp}\rho_{i}$. Similar to QG99, it assumed all turbulent energy at a given value of $k_{\perp}\rho_{i}$ resided at the parallel wavenumber governed by the extended critical balance. The H10 model assumed turbulence under nonrelativistic conditions $v_{te}/c\ll 1$ with isotropic ion and electron temperatures, strictly Alfvénic turbulence with $E_{comp}/E_{inc}=0$, balance turbulence with $Z_{0}^{+}/Z_{0}^{-}=1$, and strongly and isotropically driven turbulence with $\chi_{0}=1$ and $k_{\parallel 0}/k_{\perp 0}=1$ (the driving was therefore characterized by the isotropic driving wavenumber, $k_{0}\rho_{i}$). Furthermore, the same as QG99, the model assumed that the inertial range of the turbulence is sufficiently large that the anisotropic fluctuations predicted by the turbulent scaling relations remain at a low enough frequency relative to the ion cyclotron frequency $\omega\ll\Omega_{i}$ that the cyclotron resonance plays a negligible role; the H10 model calculated an estimate of the maximum value of $k_{0}\rho_{i}$ as a function of $(\beta_{i},T_{i}/T_{e})$ for this assumption to be satisfied. Thus, the H10 turbulent heating model predicted the damping of the turbulent cascade due to Landau damping and transit-time damping over the full range of the turbulent cascade, including MHD scales at $k_{\perp}\rho_{i}\lesssim 1$ and kinetic scales at $k_{\perp}\rho_{i}\gtrsim 1$. The H10 model predicted the partitioning of energy between ions and electrons as a function of the ion plasma beta $\beta_{i}$ and the ion-to-electron temperature ratio $T_{i}/T_{e}$, providing a simple analytical fit to the numerically calculated predictions, $Q_{i}/Q_{e}(\beta_{i},T_{i}/T_{e})$. In short, the model found that $Q_{i}/Q_{e}$ is a montonically increasing function of $\beta_{i}$ with weak dependence on $T_{i}/T_{e}$, where the transition from $Q_{i}/Q_{e}<1$ to $Q_{i}/Q_{e}>1$ happens at approximately $\beta_{i}\sim 1$.
@@
The Chandran 2011 (hereafter C11) turbulent heating model (Chandran et al., 2011) looked to incorporate the recently quantified damping rate due to nonlinear ion stochastic heating (Chandran et al., 2010; Chandran, 2010) into a model that also accounted for ion and electron Landau damping and transit-time damping. Similar to QG99 and H10, it assumed a sufficiently large inertial range that no significant cyclotron damping would occur since the anisotropic cascade would lead to turbulent fluctuations with low frequencies, $\omega\ll\Omega_{i}$. Focusing also on Alfvénic turbulence that transitions from an MHD Alfvén wave cascade at $k_{\perp}\rho_{i}\lesssim 1$ and to a kinetic Alfvén wave cascade at $k_{\perp}\rho_{i}\gtrsim 1$, the C11 model assumed that dissipation occured dominantly in two distinct wavenumber ranges: (i) at $k_{\perp}\rho_{i}\sim 1$, where collisionless damping via the Landau ($n=0$) resonance can lead to parallel energization of the ions and electrons and also where nonlinear stochastic heating can lead to perpendicular energization of the ions; and (ii) at $k_{\perp}\rho_{i}\gg 1$, where all remaining energy in the turbulent cascade is ultimately deposited with the electrons. The model used fitting forms for the Landau-resonant collisionless damping rates for the ions and electrons at $k_{\perp}\rho_{i}=1$ over $1010^{-3}<\beta_{i}<$ 10 and $51<T_{i}/T_{e}<$ 5. It also included an empirically derived formula for the ion stochastic heating rate at $k_{\perp}\rho_{i}=1$ (Chandran et al., 2010). This empirical prescription for the ion stochastic heating depended on two dimensionless constants: one governed the scaling of the stochastic heating rate relative to the turbulent frequencies; and the other, which appeared as a factor in the argument of the exponential function, effectively established a threshold amplitude below which ion stochastic heating was negligible. Thus, the C11 model calculated how the turbulent cascade rate at $k_{\perp}\rho_{i}=1$ is partitioned among ion Landau and transit-time damping, electron Landau and transit-time damping, ion stochastic heating, and nonlinear transfer to smaller scales with $k_{\perp}\rho_{i}>1$. The model provided a partitioning of the total turbulent heating rate into parallel ion heating by ion Landau and transit-time damping $Q_{\parallel,i}/Q$, perpendicular ion stochastic heating $Q_{\perp,i}/Q$, and electron heating $Q_{e}/Q$, each as a function of ion plasma beta $\beta_{i}$, the ion-to-electron temperature ratio $T_{i}/T_{e}$, and the amplitude of the turbulent fluctuations at the ion Larmor radius scale determined using the B06 scaling for the amplitude in terms of the driving scale, as given here by (19), indicating that the amplitude dependence can be characterized by $k_{0}\rho_{i}$.
@@
Motivated by the desire to understand the partitioning of dissipated turbulent energy between ions and electrons in hot collisionless accretion flows, the Rowan 2017 (hereafter R17) heating model (Rowan et al., 2017) took a fundamentally different approach from that adopted in the QG99, H10, and C11 turbulent heating models by focusing instead in the heating arising from transrelativistic magnetic reconnection. In this case, the protons are nonrelativistic, but the electrons can be ultrarelativistic. Utilizing a large suite of 2D particle-in-cell (PIC) simulations of anti-parallel reconnection, the R17 model studied the irreversible particle energization over a range of ion plasma beta $10^{-4}\lesssim\beta_{i}\leq 2$ and magnetization $0.1\leq\sigma_{w}\leq 10$, where $\sigma_{w}$ is defined as the ratio of the magnetic energy density to the enthalpy density. Starting with a Harris current sheet (Harris, 1962) of the antiparallel magnetic field (given by a $\tanh$ profile) with width $a\sim 40d_{e}$, they carefully separated the irreversible heating associated with an increase in the entropy from the reversible heating associated with adiabatic compression. From their synthesis of the results of a large suite of simulations varying different parameters, they fit a formula to their numerical results that provided to electron-to-total heating ratio $Q_{e}/Q$ as a function of the ion plasma beta $\beta_{i}$ and magnetization $\sigma_{w}$.
@@
Subsequently, using the same hybrid gyrokinetic code, Kawazura et al. (2020) explored the effect of compressible driving $E_{comp}/E_{inc}>0$ on the ion-to-electron heating ratio $Q_{i}/Q_{e}$ (hereafter K20). The turbulence was driven at MHD scales $k_{\perp}\rho_{i}\ll 1$ with an adjustable mixture of incompressible Alfvénic fluctuations and compressible slow magnetosonic fluctuations (fast magnetosonic wave modes are ordered out of gyrokinetic simulations by the perpendicular pressure balance that is maintained under the gyrokinetic approximation (Howes et al., 2006)). At $\beta_{i}\ll 1$, they found that $Q_{i}/Q_{e}=E_{comp}/E_{inc}$, confirming a previous theoretical prediction that compressive energy strictly heats ions at low ion plasma beta $\beta_{i}\ll 1$ (Schekochihin et al., 2019). Surprisingly, the K20 model also found that $Q_{i}/Q_{e}=E_{comp}/E_{inc}$ when the compressible driving dominates, $E_{comp}/E_{inc}\gg 1$. A simple fitting formula was constructed that simply added the term $E_{comp}/E_{inc}$ to the previous prescription for $Q_{i}/Q_{e}$ due to Alfvénic turbulence in the K19 model. Thus, the K20 model provided a simple analytical formula for $Q_{i}/Q_{e}(\beta_{i},T_{i}/T_{e},E_{comp}/E_{inc})$.
The application of the K20 model to understand the ion-to-electron heating ratio in accretion disks where the magnetorotational instability (MRI) (Balbus & Hawley, 1991; Hawley & Balbus, 1991; Balbus & Hawley, 1998) drives the turbulence requires a determination of the ratio of the compressible-to-incompressible driving $E_{comp}/E_{inc}$. The application of a rotating reduced MHD (RRMHD) model to simulate the collisional MRI turbulence threaded by a nearly aximuthal magnetic field in a 3D pseudo-spectral, shearing reduced MHD code (Kawazura, 2022) found a ratio of compressible-to-incompressible turbulent energy of $2\lesssim E_{comp}/E_{inc}\lesssim 2.5$ (Kawazura et al., 2022), providing a solid foundation for the application of the K20 model to astrophysical accretion disks.
@@ 5 Dependence of Dissipation Mechanisms on Fundamental Parameters
It is worthwhile noting an important subtlety regarding irreversible plasma heating in the thermodynamic picture of turbulent dissipation in weakly collisional plasmas (Howes, 2017). In the strongly collisional limit of fluid plasma turbulence, the dissipation of turbulence is fundamentally collisional on a microscopic level. For example, in an MHD plasma, the physical mechanisms of viscosity and resistivity arise in the limit of a non-negligible collisional mean free path of the plasma particles. These two dissipative mechanisms may be derived rigorously beginning with kinetic theory through a hierarchy of moment equations derived in the limit of small mean free path by the Chapman-Enskog procedure (Chapman & Cowling, 1970) for neutral gases, or an analogous procedure for plasma systems (Spitzer, 1962; Grad, 1963; Braginskii, 1965). Because viscosity and resistivity are ultimately collisional in nature, the resulting energy transfer from the turbulent fluctuations to the particles is irreversible, leading to an increase in the entropy of the system.
Under the weakly collisional conditions typical of many space and astrophysical plasmas, the collisionless removal of energy from the turbulent fluctuations and the ultimate conversion of that removed energy into plasma heat is a two-step process (Howes et al., 2006; Howes, 2008; Schekochihin et al., 2009; Navarro et al., 2016; Howes, 2017; Howes et al., 2018). First, collisionless damping mechanisms remove energy from the turbulent fluctuations, transferring that energy into nonthermal energy contained in non-Maxwellian fluctuations in the particle velocity distributions, an energy transfer that is reversible in principle. That free energy in the fluctuations of the velocity distribution about the equilibrium may be transferred to smaller scales in velocity space—e.g., via linear phase mixing by the ballistic term of the Boltzmann equation (Landau, 1946; Hammett et al., 1992; Snyder et al., 1997; Schekochihin et al., 2009) or via nonlinear phase mixing, a process recently denoted the entropy cascade (Dorland & Hammett, 1993; Schekochihin et al., 2009; Tatsuno et al., 2009; Plunk et al., 2010; Plunk & Tatsuno, 2011; Kawamori, 2013)—until it reaches sufficiently small velocity-space scales that an arbitrarily weak collisionality is sufficient to smooth out those fluctuations, providing an irreversible conversion of that energy to plasma heat and thereby increasing the entropy of the system (Howes, 2008; Schekochihin et al., 2009).
@@
A primary aim of this study is to provide a unified framework for characterizing the dependence of the turbulence and its dissipation on a common set of dimensionless parameters. The first step pursued here is to identify the fundamental dimensionless parameters upon which each proposed turbulent damping mechanism primarily depends, where the set of dimensionless parameters for the general case is presented in Table 2 and for the isotropic temperature case is in Table 3. The key parameter dependencies found here are summarized for each proposed damping mechanism in Table 4. Although, for any single particular mechanism, different choices of dimensionless parameters may be more natural (e.g., taking $\beta_{e}$ instead of $\beta_{i}$ in the case of magnetic reconnection), the particular sets of dimensionless parameters in Tables 2 and 3 are proposed here as a useful framework for describing all of the proposed turbulent damping mechanisms, thereby simplifying the task of detemining how different mechanisms compete with each other throughout the full parameter space.
@@
Table 4: Proposed turbulent damping mechanisms and their key dimensionless parameter dependencies.

### v2

@@ 4 Review of Previous Turbulent Heating Models
The Howes 2010 (hereafter H10) turbulent heating model (Howes, 2010) was based on a turbulent cascade model (Howes et al., 2008a) that extended the GS95 turbulence scaling from MHD scales at $k_{\perp}\rho_{i}\lesssim 1$ into the kinetic regime at $k_{\perp}\rho_{i}\gtrsim 1$, generalizing the conjecture of critical balance to account for the dispersive nature of kinetic Alfvén waves at these sub-ion length scales (Biskmap et al., 1999; Cho & Lazarian, 2004; Krishan & Mahajan, 2004; Shaikh & Zank, 2005; Schekochihin et al., 2009). The H10 model was based on three primary assumptions: (i) the Kolmogorov hypothesis that the energy cascade is determined by local interactions (Kolmogorov, 1941); (ii) the conjecture that turbulence maintains a state of critical balance at all scales (Goldreich & Sridhar, 1995); and (iii) the speculation that linear collisionless damping rates are applicable even in the presence of the strong nonlinear interactions that mediate the turbulent cascade to small scales. The cascade model used by H10 balanced at each perpendicular wavenumber $k_{\perp}\rho_{i}$ the nonlinear energy cascade rate given by the extended turbulent scaling theories with the linear collisionless damping rates at that value of $k_{\perp}\rho_{i}$. Similar to QG99, it assumed all turbulent energy at a given value of $k_{\perp}\rho_{i}$ resided at the parallel wavenumber governed by the extended critical balance. The H10 model assumed turbulence under nonrelativistic conditions $v_{te}/c\ll 1$ with isotropic ion and electron temperatures, strictly Alfvénic turbulence with $E_{comp}/E_{inc}=0$, balance turbulence with $Z_{0}^{+}/Z_{0}^{-}=1$, and strongly and isotropically driven turbulence with $\chi_{0}=1$ and $k_{\parallel 0}/k_{\perp 0}=1$ (the driving was therefore characterized by the isotropic driving wavenumber, $k_{0}\rho_{i}$). Furthermore, adopting the same argument as the QG99 model, the model assumed that the inertial range of the turbulence is sufficiently large that the anisotropic fluctuations predicted by the turbulent scaling relations remain at a low enough frequency relative to the ion cyclotron frequency, $\omega\ll\Omega_{i}$, that the cyclotron resonance plays a negligible role; the H10 model calculated an estimate of the maximum value of $k_{0}\rho_{i}$ as a function of $(\beta_{i},T_{i}/T_{e})$ for this assumption to be satisfied. Thus, the H10 turbulent heating model predicted the damping of the turbulent cascade due to Landau damping and transit-time damping over the full range of the turbulent cascade, including MHD scales at $k_{\perp}\rho_{i}\lesssim 1$ and kinetic scales at $k_{\perp}\rho_{i}\gtrsim 1$. The H10 model predicted the partitioning of energy between ions and electrons as a function of the ion plasma beta $\beta_{i}$ and the ion-to-electron temperature ratio $T_{i}/T_{e}$, providing a simple analytical fit to the numerically calculated predictions, $Q_{i}/Q_{e}(\beta_{i},T_{i}/T_{e})$. In short, the model found that $Q_{i}/Q_{e}$ is a monotonically increasing function of $\beta_{i}$ with weak dependence on $T_{i}/T_{e}$, where the transition from $Q_{i}/Q_{e}<1$ to $Q_{i}/Q_{e}>1$ happens at approximately $\beta_{i}\sim 1$.
@@
The Chandran 2011 (hereafter C11) turbulent heating model (Chandran et al., 2011) looked to incorporate the recently quantified damping rate due to nonlinear ion stochastic heating (Chandran et al., 2010; Chandran, 2010) into a model that also accounted for ion and electron Landau damping and transit-time damping. Similar to QG99 and H10, it assumed a sufficiently large inertial range that no significant cyclotron damping would occur since the anisotropic cascade would lead to turbulent fluctuations with low frequencies, $\omega\ll\Omega_{i}$. Focusing also on Alfvénic turbulence that transitions from an MHD Alfvén wave cascade at $k_{\perp}\rho_{i}\lesssim 1$ and to a kinetic Alfvén wave cascade at $k_{\perp}\rho_{i}\gtrsim 1$, the C11 model assumed that dissipation occurred dominantly in two distinct wavenumber ranges: (i) at $k_{\perp}\rho_{i}\sim 1$, where collisionless damping via the Landau ($n=0$) resonance can lead to parallel energization of the ions and electrons and also where nonlinear stochastic heating can lead to perpendicular energization of the ions; and (ii) at $k_{\perp}\rho_{i}\gg 1$, where all remaining energy in the turbulent cascade is ultimately deposited with the electrons. The model used fitting forms for the Landau-resonant collisionless damping rates for the ions and electrons at $k_{\perp}\rho_{i}=1$ over $1010^{-3}<\beta_{i}<$ 10 and $51<T_{i}/T_{e}<$ 5. It also included an empirically derived formula for the ion stochastic heating rate at $k_{\perp}\rho_{i}=1$ (Chandran et al., 2010). This empirical prescription for the ion stochastic heating depended on two dimensionless constants: one governed the scaling of the stochastic heating rate relative to the turbulent frequencies; and the other, which appeared as a factor in the argument of the exponential function, effectively established a threshold amplitude below which ion stochastic heating was negligible. Thus, the C11 model calculated how the turbulent cascade rate at $k_{\perp}\rho_{i}=1$ is partitioned among ion Landau and transit-time damping, electron Landau and transit-time damping, ion stochastic heating, and nonlinear transfer to smaller scales with $k_{\perp}\rho_{i}>1$. The model provided a partitioning of the total turbulent heating rate into parallel ion heating by ion Landau and transit-time damping $Q_{\parallel,i}/Q$, perpendicular ion stochastic heating $Q_{\perp,i}/Q$, and electron heating $Q_{e}/Q$, each as a function of ion plasma beta $\beta_{i}$, the ion-to-electron temperature ratio $T_{i}/T_{e}$, and the amplitude of the turbulent fluctuations at the ion Larmor radius scale determined using the B06 scaling for the amplitude in terms of the driving scale, as given here by (19), indicating that the amplitude dependence can be characterized by $k_{0}\rho_{i}$.
@@
Motivated by the desire to understand the partitioning of dissipated turbulent energy between ions and electrons in hot collisionless accretion flows, the Rowan 2017 (hereafter R17) heating model (Rowan et al., 2017) took a fundamentally different approach from that adopted in the QG99, H10, and C11 turbulent heating models by focusing instead on the heating arising from transrelativistic magnetic reconnection. In this case, the protons are nonrelativistic, but the electrons can be ultrarelativistic. Utilizing a large suite of 2D particle-in-cell (PIC) simulations of anti-parallel reconnection, the R17 model studied the irreversible particle energization over a range of ion plasma beta $10^{-4}\lesssim\beta_{i}\leq 2$ and magnetization $0.1\leq\sigma_{w}\leq 10$, where $\sigma_{w}$ is defined as the ratio of the magnetic energy density to the enthalpy density. Starting with a Harris current sheet (Harris, 1962) of the antiparallel magnetic field (given by a $\tanh$ profile) with width $a\sim 40d_{e}$, they carefully separated the irreversible heating associated with an increase in the entropy from the reversible heating associated with adiabatic compression. From their synthesis of the results of a large suite of simulations varying different parameters, they fit a formula to their numerical results that provided to electron-to-total heating ratio $Q_{e}/Q$ as a function of the ion plasma beta $\beta_{i}$ and magnetization $\sigma_{w}$.
@@
The first of the two disagreements above between the H10 model and the K19 model—where the numerical simulations found an effective ceiling at $Q_{i}/Q_{e}\simeq 30$ at $\beta_{i}\gg 1$—was recently addressed by Gorman & Klein (2024). They examined in detail the small range of perpendicular wavenumber scales around $k_{\perp}\rho_{i}\sim 1$ where the linear dispersion relation predicts that the Alfvén wave solution becomes non-propagating at $\beta_{i}\gg 1$. For turbulent cascade models assuming critical balance and strictly local (in scale space) interactions mediating the turbulent cascade, this gap in the Alfvén wave propagation leads to the turbulent cascade being terminated at the onset of the gap with the bulk of the turbulent energy transferred to the ions, leading to the potential overestimation of $Q_{i}/Q_{e}$ in the $\beta_{i}\gg 1$ limit. Gorman & Klein (2024) (hereafter GK24) tested the local cascade model underlying the H10 model against a more refined cascade model, the weakened cascade model (Howes et al., 2011). The weakened cascade model is a nonlocal cascade model that accounts both for the weakening of the local nonlinear interactions as damping taps the turbulent energy and for nonlocal interactions by large-scale shearing and small-scale diffusion. It was found that the inclusion of nonlocal interactions to the nonlinear energy transfer prevents the termination of the cascade at the ion scales, leading to predictions for $Q_{i}/Q_{e}$ at $\beta_{i}\gg 1$ (Gorman & Klein, 2024) that agree much more closely with the results of the suite of simulations by Kawazura et al. (2019): the resulting GK24 model produced by this study includes updated coefficients for the H10 formula for $Q_{i}/Q_{e}(\beta_{i},T_{i}/T_{e})$.
Relaxing the limitation of their simulations to driving by strictly Alfvénic turbulent fluctuations, Kawazura et al. (2020) used the same hybrid gyrokinetic code (Kawazura & Barnes, 2018) to explore the effect of compressible driving $E_{comp}/E_{inc}>0$ on the ion-to-electron heating ratio $Q_{i}/Q_{e}$ (hereafter K20). The turbulence was driven at MHD scales $k_{\perp}\rho_{i}\ll 1$ with an adjustable mixture of incompressible Alfvénic fluctuations and compressible slow magnetosonic fluctuations [fast magnetosonic wave modes are ordered out of gyrokinetic simulations by the perpendicular pressure balance that is maintained under the gyrokinetic approximation (Howes et al., 2006)]. At $\beta_{i}\ll 1$, they found that $Q_{i}/Q_{e}=E_{comp}/E_{inc}$, confirming a previous theoretical prediction that compressive energy strictly heats ions at low ion plasma beta $\beta_{i}\ll 1$ (Schekochihin et al., 2019). Surprisingly, the K20 model also found that $Q_{i}/Q_{e}=E_{comp}/E_{inc}$ when the compressible driving dominates, $E_{comp}/E_{inc}\gg 1$. A simple fitting formula was constructed that simply added the term $E_{comp}/E_{inc}$ to the previous prescription for $Q_{i}/Q_{e}$ due to Alfvénic turbulence in the K19 model. Thus, the K20 model provided a simple analytical formula for $Q_{i}/Q_{e}(\beta_{i},T_{i}/T_{e},E_{comp}/E_{inc})$.
The application of the K20 model to understand the ion-to-electron heating ratio in accretion disks where the magnetorotational instability (MRI) (Balbus & Hawley, 1991; Hawley & Balbus, 1991; Balbus & Hawley, 1998) drives the turbulence requires a determination of the ratio of the compressible-to-incompressible driving $E_{comp}/E_{inc}$. The application of a rotating reduced MHD (RRMHD) model to simulate the collisional MRI turbulence threaded by a nearly azimuthal magnetic field in a 3D pseudo-spectral, shearing reduced MHD code (Kawazura, 2022) found a ratio of compressible-to-incompressible turbulent energy of $2\lesssim E_{comp}/E_{inc}\lesssim 2.5$ (Kawazura et al., 2022), providing a solid foundation for the application of the K20 model to astrophysical accretion disks.
@@ 5 Dependence of Dissipation Mechanisms on Fundamental Parameters
It is worthwhile first noting an important subtlety regarding irreversible plasma heating in the thermodynamic picture of turbulent dissipation in weakly collisional plasmas (Howes, 2017). In the strongly collisional limit of fluid plasma turbulence, the dissipation of turbulence is fundamentally collisional on a microscopic level. For example, in an MHD plasma, the physical mechanisms of viscosity and resistivity arise in the limit of a non-negligible collisional mean free path of the plasma particles. These two dissipative mechanisms may be derived rigorously beginning with kinetic theory through a hierarchy of moment equations derived in the limit of small mean free path by the Chapman-Enskog procedure (Chapman & Cowling, 1970) for neutral gases, or an analogous procedure for plasma systems (Spitzer, 1962; Grad, 1963; Braginskii, 1965). Because viscosity and resistivity are ultimately collisional in nature, the resulting energy transfer from the turbulent fluctuations to the particles is irreversible, leading to an increase in the entropy of the system.
Under the weakly collisional conditions typical of many space and astrophysical plasmas, the collisionless removal of energy from the turbulent fluctuations and the ultimate conversion of that removed energy into plasma heat is a two-step process (Howes et al., 2006; Howes, 2008; Schekochihin et al., 2009; Navarro et al., 2016; Howes, 2017; Howes et al., 2018). First, collisionless damping mechanisms remove energy from the turbulent fluctuations, transferring that energy into nonthermal internal energy contained in non-Maxwellian fluctuations of the particle velocity distributions, an energy transfer that is reversible in principle. That free energy in the fluctuations of the velocity distribution about the equilibrium may be transferred to smaller scales in velocity space—e.g., via linear phase mixing by the ballistic term of the Boltzmann equation (Landau, 1946; Hammett et al., 1992; Snyder et al., 1997; Schekochihin et al., 2009) or via nonlinear phase mixing, a process often denoted the entropy cascade (Dorland & Hammett, 1993; Schekochihin et al., 2009; Tatsuno et al., 2009; Plunk et al., 2010; Plunk & Tatsuno, 2011; Kawamori, 2013)—until it reaches sufficiently small velocity-space scales that an arbitrarily weak collisionality is sufficient to smooth out those fluctuations, providing an irreversible conversion of that energy to plasma heat and thereby increasing the entropy of the system (Howes et al., 2006; Howes, 2008; Schekochihin et al., 2009).
@@
A primary aim of this study is to provide a unified framework for characterizing the dependence of the turbulence and its dissipation on a common set of dimensionless parameters. The first step pursued here is to identify the fundamental dimensionless parameters upon which each proposed turbulent damping mechanism primarily depends, where the set of dimensionless parameters for the general case is presented in Table 3 and for the isotropic temperature case in Table 4. The key parameter dependencies for each proposed damping mechanism found here are summarized in Table 5. Although, for any single particular mechanism, different choices of dimensionless parameters may be more natural (e.g., taking $\beta_{e}$ instead of $\beta_{i}/\tau$ in the case of magnetic reconnection), the particular sets of dimensionless parameters in Tables 3 and 4 are proposed here as a useful framework for describing all of the proposed turbulent damping mechanisms, thereby simplifying the task of determining how different mechanisms compete with each other throughout the full parameter space.
@@
Table 5: Proposed turbulent damping mechanisms and their key dimensionless parameter dependencies for the case of a fully ionized hydrogenic plasma. The ion-to-electron mass ratio $\mu$ is listed in parentheses since it is not a physically variable parameter, but it is often adjusted in numerical studies for computational efficiency.